A bias-neutral venture capitalist is contemplating investing $3.5 million in a start-up company. Company net income at end of year four is expected to be $3 million. Given the four year length of his investment, his firm requires a 50% internal rate of return compounded annually. Given that the start-up company’s product line is a novel computer technology, rapid growth is expected and therefore the price to earnings ratio is expected to be 55. Neither future investments nor change in earnings projection is expected prior to the end of the investment term. There are currently 3 million shares outstanding. Using the Sahlman Ownership proportion Formula, compute rounded to the nearest thousandth the percentage of ownership in a start-up company that the venture capitalists should seek.
Answer #1

The VC must realise a 50% return on investment after 4 years.

Therefore,

His investment value after 4 years = Investment*(1+r)^{n}

Investment worth after 4 years = 3.5*(1+0.5)^{4} = $17.71875 Mn

Net Income(year 4) = $3Mn

Expected Price/ Earning = 55

Value of company at end of year 4 = Net Income * P/E ratio = 55*$3Mn = $165Mn

For VC to receive $17.71Mn at end of year 4 out of $165Mn Company value,

Ownership% required = 17.71875/165 = 10.7386%

Please rate a thumbs up. Feel free to comment for any query.